# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ 5 Discussion on the range of $\gamma$
In the final section, we give a conjecture on the range of $\gamma$. Consider the rough of singular Burgers equation driven by $|D|^{\beta}\xi,$
$u_{t}-\Lambda^{\gamma}u=\frac{1}{2}\partial_{x}(u^{2})+|D|^{\beta}\xi.$ (5.1)
where $\beta\geq 0$. Fix $\zeta=|D|^{\beta}\xi$, For dyadic number $N\neq 0$, fix $\zeta^{N}:=\mathcal{F}^{-1}(1_{|k|\sim N}\mathcal{F}\xi(k))$, which means the frequency of $\xi$ is focus on $\frac{1}{2}N\leq|k|\leq N$. Then the Fourier mode of the linear evolution $X:=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\zeta^{N}\mathrm{d}s$ can be written as
@@
Its not difficult to calculate that $G_{k}(t)$ from a collection of independent Gaussian variable with $\mathbb{E}[G_{k}(t)]\sim 1$, then we can get $X$ belong to $C_{T}H^{\frac{\gamma}{2}-\beta-\frac{1}{2}}$. Now we calculate the "first nonlinear iteration", fix $u^{(1)}(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(X^{2})\mathrm{d}s$, and fix $u_{k}^{(1)}(t)$ is the Fourier mode
@@
We call $(\beta,\gamma)$ is a "probability admissible pair" for Burgers equation, if (5.2) satisfied. We can simply get that $(0,\gamma)$ for $\gamma>\frac{4}{3}$ and $(1,2)$ are probability admissible pairs, which the first range was studied in our paper and the second in [GP17].
@@ 5.1 Future directions
In future work, we will continue our investigation of stochastic Burgers equation, but in a more singular case. For example, $\gamma=2$ and $\beta=1$ for (5.1). Note that $u$ gains the regularity as $X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$ has, by Lemma 2.7, we priori get $u\in-\frac{1}{2}-\delta$ for any $\delta>0$. It shows that $\partial_{x}(u^{2})$ is maybe ill-defined since the condition of Lemma 2.1 is not hold! To solve this, a common approach is to study its generalized solution. More preciesly, let $u^{\epsilon}$ satisfied
$\partial_{t}u^{\epsilon}-\Lambda^{\gamma}u^{\epsilon}=\frac{1}{2}D(u^{\epsilon})^{2}+|D|^{\beta}\xi_{\epsilon},\quad u^{\epsilon}(0)=u^{\epsilon}_{0}+Y_{\epsilon}(0),$ (5.3)
and prove the $u_{\epsilon}$ has a limit $u$ in probability sense. This idea was first proposed by Martin Hairer and applied to the KPZ equation in [Hai13]. In his subsequent work [Hai14], they established a comprehensive theory of regularity structures, which resolved the problem of defining solutions for a broad class of singular SPDEs, earning him the Fields Medal. There, they build a equation in modeled distribution space $\mathcal{D}^{\gamma,\eta}_{P}$ on $[0,T]$ such that
@@
Our next target is to build a generalized equation in the classical space for example $W^{\alpha}$, and prove that the generalized solution of singular Burgers equation 5.1 can be back to the generalized equation.
@@ Appendix A Estimate for heat flow
In this chapter, we give some estimate for heat flow. Define operator $P(t)$ as
$\mathcal{F}(P(t)f)=e^{-t|\xi|^{\gamma}}\mathcal{F}f.$ (A.1)
for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Its not difficult to see that $P(t)f$ is the solution of
$\partial_{t}u-\Delta u=0,\quad u(0)=f.$ (A.2)
For the operator, we have following estimate
@@
For $\gamma>0$, we have the estimate
@@
if the support of $\hat{f}$ is belong to some $\lambda\mathcal{C}$, where $\mathcal{C}$ is a annual.
@@
The proof is similar as Lemma 2.4 in [BCD11], which only need to notice that $\partial_{x}^{\beta}e^{-t|\xi|^{\gamma}}\leq C(1+t)^{|\beta|}e^{-ct}$ for $\xi\in\lambda\mathcal{C}$ will not cause difficult when $\gamma>0$, since $\xi$ take away from 0. ∎
Following this useful Lemma, we can get
@@
Let $p,q\in[1,\infty]$, $\beta\in\mathbb{Z}_{+}$, then for $\gamma\geq 1$, we have
$\|\partial_{x}^{\beta}P(t)f\|_{L^{p}}\lesssim t^{-\frac{\beta}{\gamma}-\frac{1}{\gamma}(\frac{1}{r}-\frac{1}{p})}\|f\|_{L^{r}}.$
@@
We rewrite
$\partial_{x}^{\beta}P(t)f=G(t,x)\ast f,$
where $g(t,x)=(\mathcal{F}^{-1}(ik)^{\beta}e^{-t|k|^{\gamma}})$. By Young inequality, it suffice to show that for $\frac{1}{q}=1+\frac{1}{p}-\frac{1}{r}$, we have $\|G(t,x)\|_{L^{q}}\leq t^{-\frac{\beta}{\gamma}-(1-\frac{1}{q})\frac{1}{\gamma}}$. In fact, rewrite $G(t,x)$ as
$G(t,x)=t^{-\frac{\gamma}{\beta}}\mathcal{F}^{-1}(it^{\frac{1}{\gamma}}k)^{\beta}e^{-|t^{\frac{1}{\gamma}}k|^{\gamma}}.$
Since $g(k)=k^{\beta}e^{-|k|^{\gamma}}$ is belong to $\mathcal{S}(\mathbb{R})$, we get $\check{g}(x)\in\mathcal{S}(\mathbb{R})\subset L^{q}(\mathbb{R})$. By Lemma 1.24 in [BCD11], the bound
$\|G(t,x)\|_{L^{q}(\mathbb{T})}\lesssim t^{-\frac{\beta}{\gamma}-\frac{1}{\gamma}}\|\check{g}(t^{-\frac{1}{\gamma}}x)\|_{L^{q}(\mathbb{T})}\lesssim t^{-\frac{\beta}{\gamma}-(1-\frac{1}{q})\frac{1}{\gamma}}\|\check{g}(\cdot)\|_{L^{q}(\mathbb{R})}\leq t^{-\frac{\beta}{\gamma}-(1-\frac{1}{q})\frac{1}{\gamma}},$
which we prove our claim. ∎
For Besov space $B_{p,r}^{s}$, we have result
Lemma A.3.
Let $1\leq p,r\leq\infty$, $\alpha\in\mathbb{R}$ and $\gamma\geq 1$, then we have following estimate
$\|P(t)f\|_{B_{p,r}^{\alpha}}\lesssim\|f\|_{B_{p,r}^{\alpha}}.$ (A.3)
Let $\delta\geq 0$ and $t>0$, we have
$\|P(t)f\|_{B_{p,r}^{\alpha+\delta}}\leq t^{-\frac{\delta}{\gamma}}\|f\|_{B_{p,r}^{\alpha}}.$
$\|f\|_{B_{p,r}^{\alpha+\delta}}\lesssim\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{B_{p,r}^{\alpha}}.$
Proof.
For $j\geq 0$, by Lemma A.1, for $t>0$, we have
@@
Multiplying $2^{j\alpha}$ both sides and taking the $l^{r}$-norm with respect to $1\leq j$, we prove the second result.
For the last claim, by the last result
$\|f\|_{B_{p,r}^{\alpha+\delta}}\leq$ $\|P(t)f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|P(t-s)\mathcal{L}f(s)\|_{B_{p,r}^{\alpha+\delta}}$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|(t-s)^{-\frac{\delta}{\gamma}}\mathcal{L}f(s)\|_{B_{p,r}^{\alpha}}$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f(s)\|_{B_{p,r}^{\alpha}}$
@@
The similar result is also valid for $C_{T}^{\alpha}L^{p}$. We have the following lemma
@@
Let $\gamma\geq 1$ and $\alpha\in\mathbb{R}_{+}$, for any $T\geq 0$ we have estimate
@@
Moreover, let $\alpha\in(0,\gamma)$, for $\delta\in[0,\gamma)$ satisfied $0<\alpha+\delta<\gamma$ and $0\leq T\leq 1$, it has the estimate
@@
$\|f(t)-f(s)\|_{L^{2}}\leq\int_{s}^{t}\|P(t-\tau)\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau+\int_{0}^{s}\|[P(t-\tau)-P(s-\tau)]\mathcal{L}f(\tau)\|\mathrm{d}\tau.$
@@
For the second term, we estimate as
@@
Summary all, from the $C_{T}^{\frac{\alpha+\delta}{\gamma}}L^{2}$, we have
@@
To clear that, fix $\varphi(t,z)=\frac{e^{-|z|^{\gamma}}}{1+t^{-\frac{2}{\gamma}}|z|^{2}}.$ We have $\tilde{P}(t)f=\varphi(t,t^{\frac{1}{\gamma}}D)f=\tilde{G}(t,x)f$. Fixing $t\in[0,T]$, its easy to see that $\mathcal{F}^{-1}\varphi(t,z)\in L^{1}(\mathbb{R})$. In fact, by Young’s inequality
@@
Since $cz^{2}>0$, it’s obviously that $v,v^{\prime},v^{\prime\prime}\leq 1$ which shows that
@@
On the other hand, since all partial derivatives of $u=e^{-|z|^{\gamma}}$ is decay faster than rational function for $\gamma\geq 1$, we have
@@
$\lesssim$ $c^{-1}$
Note that $c=t^{-\frac{2}{\gamma}}$ and $d=1$. Then there exists $j_{0}$ make sure that $2^{j_{0}}t^{\frac{1}{\gamma}}\geq 1$ such that
@@
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}$
For $1\leq j<j_{0}$, we have the truth
@@
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f\|_{L^{p}}$
Summary above, we prove following lemma
@@
Let $\gamma\geq 1$, $\tilde{P}(t)$ is defined as $\mathcal{F}(\tilde{P}(t)(f))=\frac{e^{-t|\xi|^{\gamma}}}{1+|\xi|^{2}}\mathcal{F}f$. Fix some $T>0$, then for any $t\in(0,T]$, $\alpha\in\mathbb{R}$, $\delta\in(0,\gamma)$, and $u\in S^{\prime}(\mathbb{R})$, we have
@@
Moreover, let $\tilde{L}f=(\partial_{t}-\Lambda^{\gamma})(1-\partial_{x}^{2})^{-1}f\in B_{p,r}^{\alpha}$, then for $\delta\in[0,\gamma+2)$ and $t\leq 1$, we have estimate
@@ Appendix B Commutator estimates
Fix $f(x)=|x|^{\gamma}$, then we have $D(u\prec v)-u\prec Dv\in W_{p}^{\alpha+\beta-\gamma}$ for $u\in\mathcal{C}^{\alpha}$, $\alpha\in(0,1)$ and $v\in W_{p}^{\beta}$.
Following Lemma is proved
@@
Assume that $\alpha\in(0,1)$, $\beta,\gamma\in\mathbb{R}$ and satisfied $\beta+\gamma<0$, $\alpha+\beta+\gamma>0$, define $C(f,g,h)=(f\prec g)\circ h-f(g\circ h)$, then there is estimate that
@@
Its suffice to estimate above two terms. For any fixed $k$, we have estimate
@@
Similar as above estimate, we finish our proof. ∎
@@
Let $\alpha\in(0,\gamma)$ and $\beta\in\mathbb{R}$. Then
$\|f\prec\!\!\!\prec g-f\prec g\|_{C_{T}W^{\alpha+\beta}}\lesssim\|f\|_{\mathcal{W}_{\infty}^{\alpha}(T)}\|g\|_{C_{T}W^{\beta}}.$ (B.1)
Moreover, define operator $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then for $\gamma\geq 1$, it has
$\|\mathcal{L}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec(\mathcal{L}g)\|_{C_{T}W^{\alpha+\beta-\gamma}}\lesssim\|f\|_{C_{T}\mathcal{C}^{\alpha}}\|g(t)\|_{C_{T}W^{\beta}}.$ (B.2)
@@
We only proof the case of $H^{\alpha}$, the case of $\mathcal{C}^{\alpha}$ can be similar. Firstly, control
@@
Multiplying $2^{-j(\alpha+\beta)}$ both sides to the above inequality and sum over $l^{2}$, we prove (B.1).
For the second result, it’s trivial for fractional Laplacian commutator by Bernstein inequality and Lemma B.1. We only estimate

### v2

@@ 5 Discussion on the range of $\gamma$
In the final section, we propose a conjecture regarding the admissible range of the dissipation parameter $\gamma$ for the rough or singular Burgers equation driven by $|D|^{\beta}\xi.$
$u_{t}-\Lambda^{\gamma}u=\partial_{x}(u^{2})+|D|^{\beta}\xi.$ (5.1)
where $\beta\geq 0$. For $\zeta=|D|^{\beta}\xi$, we fix a dyadic number $N\neq 0$, and set $\zeta^{N}:=\mathcal{F}^{-1}(1_{|k|\sim N}\mathcal{F}\xi(k))$, which means the frequency of $\xi$ is focus on $\frac{1}{2}N\leq|k|\leq N$. Then the Fourier mode of the linear evolution $X:=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\zeta^{N}\mathrm{d}s$ can be written as
@@
It’s not difficult to calculate that $G_{k}(t)$ from a collection of independent Gaussian variable with $\mathbb{E}[G_{k}(t)]\sim 1$, then we can get $X$ belong to $C_{T}H^{\frac{\gamma}{2}-\beta-\frac{1}{2}}$. Now we calculate the "first nonlinear iteration". For $u^{(1)}(t):=\int_{0}^{t}e^{(t-s)\Lambda^{\gamma}}\partial_{x}(X^{2})\mathrm{d}s$, and let $u_{k}^{(1)}(t)$ be its Fourier mode
@@
We call a pair $(\beta,\gamma)$ “probabilistic admissible pair" for the Rough or singular Burgers equation on $H^{s}$ space(In the $\mathcal{C}^{s}$ space, it will be different), if condition (5.2) holds. It is immediate to verify that the pair $(0,\gamma)$ for $\gamma>\frac{4}{3}$ and $(\frac{1}{2},\frac{3}{2})$ are probabilistic admissible pairs; the former regime is treated in the present paper, while the latter will be our future directions.
@@ 5.1 Future directions
In future work, we intend to extend our analysis to more singular regimes of the stochastic Burgers equation. A primary example is the parameter choice $\gamma\geq\frac{3}{2}$ and $\beta=\frac{1}{2}$ in (5.1). The solution $u$ inherits the spatial regularity of the stochastic convolution $X(t)=\int_{0}^{t}P(t-s)\xi(s)\mathrm{d}s$. According to the Lemma 2.7, we have a priori $u\in W^{-\frac{1}{2}-\delta}$ for any $\delta>0$. At such low regularity, the nonlinear term $\partial_{x}(u^{2})$ may become ill-defined, as the assumptions required for Lemma 2.1 no longer hold. A standard way to circumvent this difficulty is to work with generalized solutions. More precisely, we consider a regularized solution $u^{\epsilon}$ satisfying
$\partial_{t}u^{\epsilon}-\Lambda^{\gamma}u^{\epsilon}=D(u^{\epsilon})^{2}+|D|^{\beta}\xi_{\epsilon},\quad u^{\epsilon}(0)=u^{\epsilon}_{0}+Y_{\epsilon}(0),$ (5.3)
and prove that $u^{\epsilon}$ converges in probability to a limiting process $u$. This approach was originally introduced by Martin Hairer in his analysis of the KPZ equation [Hai13]. In his subsequent foundational work [Hai14], he developed the comprehensive framework of regularity structures, which provides a systematic way to define solutions for a broad class of singular stochastic partial differential equations—contributions that were recognized with the Fields Medal. Within that framework, the solution is constructed in a space of modelled distributions $\mathcal{D}^{\gamma,\eta}_{P}$ over $[0,T]$ and satisfies the fixed-point equation
@@ Appendix A Estimate for heat flow
In this chapter, we give some estimates for heat flow. Defining operator $P(t)$ as
$\mathcal{F}(P(t)f)=e^{-t|\xi|^{\gamma}}\mathcal{F}f,$ (A.1)
for $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$, it’s not difficult to see that $P(t)f$ is the solution of
$\partial_{t}P(t)f-\Lambda^{\gamma}P(t)f=0,\quad u(0)=f.$ (A.2)
For the operator, we firstly claim that it’s the linear bounded operator in $L^{p}$ for $p\geq 1$. In fact, setting $\varphi(z)=e^{-|z|^{\gamma}}$, then $P(t)$ can be written as $\varphi(t^{\frac{1}{\gamma}}D)$. Noting that $\varphi\in L^{1}$, it suffice to show $\mathcal{F}\varphi\in L^{1}$, then for any $t\geq 0$, we have
$\|P(t)f\|_{L^{p}}=\|\mathcal{F}^{-1}\varphi(t^{\frac{1}{\gamma}}\cdot)\ast f\|_{L^{p}}\lesssim\|f\|_{L^{p}}.$ (A.3)
We can easily get $\mathcal{F}\varphi$ belongs to $L^{1}$ since its the density of a symmetric $\gamma$-stable random variable for $\gamma\in(0,2]$.
@@
For $\gamma>0$, $\lambda\in\mathbb{Z}/\{0\}$, we have the estimate
@@
if the support of $\hat{f}$ is belong to some $\lambda\mathcal{C}$, where $\mathcal{C}$ is a fixed annual.
@@
The proof is similar as Lemma 2.4 in [BCD11], which only need to notice that $\partial_{x}^{\beta}e^{-t|\xi|^{\gamma}}\leq C(1+t)^{|\beta|}e^{-ct}$ for $\xi\in\lambda\mathcal{C}$ will not cause difficult when $\gamma>0$, since $|\xi|$ has a uniformly lower bound for $\lambda\in\mathbb{Z}/\{0\}$. ∎
For Besov space $B_{p,r}^{s}$, we have the following result
@@
Let $1\leq p,r\leq\infty$, $\alpha\in\mathbb{R}$ and $\gamma>0$, then for $t\geq 0$, we have the following estimate
$\|P(t)f\|_{B_{p,r}^{\alpha}}\lesssim\|f\|_{B_{p,r}^{\alpha}}.$ (A.4)
Let $\delta\geq 0$ and $0<t\leq 1$, we have
$\|P(t)f\|_{B_{p,r}^{\alpha+\delta}}\leq t^{-\frac{\delta}{\gamma}}\|f\|_{B_{p,r}^{\alpha}}.$ (A.5)
$\|f\|_{B_{p,r}^{\alpha+\delta}}\lesssim\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{t}B_{p,r}^{\alpha}}.$ (A.6)
@@
For fixed $j\geq 0$, by Lemma A.1, for $t>0$, we have
@@
Multiplying $2^{j(\alpha+\delta)}$ both sides and taking the $l^{r}$-norm with respect to $1\leq j$, we prove the second result.
For (A.6), by (A.4) and (A.5)
$\|f\|_{B_{p,r}^{\alpha+\delta}}\leq$ $\|P(t)f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|P(t-s)\mathcal{L}f(s)\|_{B_{p,r}^{\alpha+\delta}}\mathrm{d}s$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+\int_{0}^{t}\|(t-s)^{-\frac{\delta}{\gamma}}\mathcal{L}f(s)\|_{B_{p,r}^{\alpha}}\mathrm{d}s$
$\lesssim$ $\|f(0)\|_{B_{p,r}^{\alpha+\delta}}+t^{1-\frac{\delta}{\gamma}}\|\mathcal{L}f\|_{C_{t}B_{p,r}^{\alpha}}$
@@
A similar result carries over to the space $C_{T}^{\alpha}L^{2}$ and $C_{T}^{\alpha}L^{\infty}$, as stated in the following lemma.
@@
Let $\gamma>0$ and $\alpha\in\mathbb{R}_{+}$, for any $T\geq 0$ we have the following estimate
@@
Moreover, let $\alpha\in(0,\gamma)$, $\delta\in[0,\gamma)$ satisfy $0<\alpha+\delta<\gamma$ and $0\leq T\leq 1$, it has the estimate
@@
$\|f(t)-f(s)\|_{L^{2}}\leq\int_{s}^{t}\|P(t-\tau)\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau+\int_{0}^{s}\|[P(t-\tau)-P(s-\tau)]\mathcal{L}f(\tau)\|_{L^{2}}\mathrm{d}\tau.$
@@
For the second term, we have estimate
@@
Combining the above estimate, by the definition of $C_{T}^{\frac{\alpha+\delta}{\gamma}}L^{2}$, we have
@@
To clear that, fix $\varphi(t,z)=\frac{e^{-|z|^{\gamma}}}{1+t^{-\frac{2}{\gamma}}|z|^{2}}.$ We have $\tilde{P}(t)f=\varphi(t,t^{\frac{1}{\gamma}}D)f=\tilde{G}(t,x)f$. For fixed $t\in[0,T]$, it’s easy to see that $\mathcal{F}^{-1}\varphi(t,z)\in L^{1}(\mathbb{R})$. In fact, by Young’s inequality and (A.1)
@@
Since $cz^{2}>0$, it’s obviously that $|v|,|v^{\prime}|,|v^{\prime\prime}|\lesssim 1$ which shows that
@@
On the other hand, since all partial derivatives of $u=e^{-|z|^{\gamma}}$ is decay faster than rational function for $\gamma>0$, we have
@@
$\lesssim$ $c^{-1}.$
Let $c=t^{-\frac{2}{\gamma}}$ and $d=1$, then there exists $j_{0}$ satisfying that $2^{j_{0}}t^{\frac{1}{\gamma}}\geq 1$ such that
@@
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f(t)\|_{L^{p}}.$
For $1\leq j<j_{0}$, we have the estimate
@@
$\lesssim$ $\min(t^{\frac{2}{\gamma}},1)t^{-\frac{\delta}{\gamma}}2^{-j\delta}\|\Delta_{j}f(t)\|_{L^{p}}.$
Combining the proceeding results, we obtain the following lemma
@@
Let $\gamma>0$, $\tilde{P}(t)$ satisfy $\mathcal{F}(\tilde{P}(t)(f))=\frac{e^{-t|\xi|^{\gamma}}}{1+|\xi|^{2}}\mathcal{F}f$. Fixing some $T>0$, then for any $t\in(0,T]$, $\alpha\in\mathbb{R}$, $\delta\in(0,\gamma)$, and $u\in S^{\prime}(\mathbb{R})$, we have the following estimate
@@
Moreover, let $\tilde{L}f=(\partial_{t}-\Lambda^{\gamma})(1-\partial_{x}^{2})^{-1}f\in B_{p,r}^{\alpha}$, then for $\delta\in[0,\gamma+2)$ and $t\leq 1$, we have the following estimate
@@ Appendix B Commutator estimates
Fixing $f(x)=|x|^{\gamma}$, then we have $\Lambda^{\gamma}(u\prec v)-u\prec\Lambda^{\gamma}v\in W_{p}^{\alpha+\beta-\gamma}$ for $u\in\mathcal{C}^{\alpha}$, $\alpha\in(0,1)$ and $v\in W_{p}^{\beta}$ for $\alpha\in(0,1)$.
@@
Assume that $\alpha\in(0,1)$, $\beta,\gamma\in\mathbb{R}$ and satisfies $\beta+\gamma<0$, $\alpha+\beta+\gamma>0$, let $C(f,g,h)=(f\prec g)\circ h-f(g\circ h)$, then we have the following estimate
@@
It’s suffice to estimate above two terms. For any fixed $k$, we have estimate
@@
Combining all estimates above, we finish our proof. ∎
@@
Let $\alpha\in(0,\gamma)$ and $\beta\in\mathbb{R}$. Then for $t>0$, we have the following estimate
$\|f\prec\!\!\!\prec g-f\prec g\|_{W^{\alpha+\beta}}\lesssim\|f\|_{\mathcal{W}_{\infty}^{\alpha}(t)}\|g\|_{W^{\beta}}.$ (B.1)
Moreover, define operator $\mathcal{L}=\partial_{t}-\Lambda^{\gamma}$, then for $\gamma>0$ and $\alpha\in(0,1)$, we have the following estimate
$\|\mathcal{L}(f\prec\!\!\!\prec g)(t)-(f\prec\!\!\!\prec\mathcal{L}g)(t)\|_{W^{\alpha+\beta-\gamma}}\lesssim\|f\|_{\mathcal{W}_{\infty}^{\alpha}(t)}\|g\|_{W^{\beta}}.$ (B.2)
@@
We only proof the case of $H^{\alpha}$, the case of $\mathcal{C}^{\alpha}$ is similar. Firstly, control
@@
Multiplying $2^{-j(\alpha+\beta)}$ both sides to the above inequality and summing over $j$ in the $l^{2}$, we prove (B.1).
For the second result, we firstly control the term of fractional Laplacian commutator
$\|\Lambda^{\gamma}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec\Lambda^{\gamma}g\|_{C_{T}W^{\alpha+\beta-\gamma}}$
For $f\in C_{t}\mathcal{C}^{\alpha}$, rewriting $\Lambda^{\gamma}(f\prec\!\!\!\prec g)(t)-f\prec\!\!\!\prec\Lambda^{\gamma}g(t)$ as
$\sum_{j\geq 1}\int_{0}^{t}2^{\gamma j}$ $\varphi(2^{\gamma j}(t-s))[\Lambda^{\gamma}(S_{j-1}f(s)\Delta_{j}g(t))-S_{j-1}f(s)\Lambda^{\gamma}(\Delta_{j}g(t))]\mathrm{d}s$
$=$ $\sum_{j\geq 1}\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\mathrm{d}s$
where $\tilde{\rho}$ is a smooth function supported in an annuals and with value 1 on a neighborhood of $\mathrm{Supp}\rho+\mathrm{Supp}\chi(\cdot/4)$. Observing that there exists $N_{0}$ such that
$\forall j\geq N_{0},\quad\Lambda^{\gamma}\widetilde{\Delta}_{j}=2^{j\gamma}(|\cdot|^{\gamma}\tilde{\rho})(2^{-j}D).$
Then by Lemma 2.97 in [BCD11], we have for any $j\geq N_{0}$
$\|[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\|_{L^{p}}\leq C2^{j(\gamma-1)}\|S_{j-1}f(s)\|_{L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}},$
and for $1\leq j<N_{0}$, we have
$\|[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\|_{L^{p}}\leq C2^{j(\gamma-1)}\|\nabla S_{j-1}f(s)\|_{L^{\infty}}\|\Delta_{j}g(t)\|_{L^{p}}$
Since $\|\nabla S_{j-1}f(s)\|_{L^{\infty}}\leq C2^{j(1-\alpha)}\|\nabla f(s)\|_{\mathcal{C}^{\alpha-1}}\leq C2^{j(1-\alpha)}\|f\|_{C_{t}\mathcal{C}^{\alpha}}$ for $\alpha<1$ and $0<s<t$. then we have
$2^{j(\beta+\alpha-\gamma)}\|\int_{0}^{t}2^{\gamma j}\varphi(2^{\gamma j}(t-s))[S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]\Delta_{j}g(t)\mathrm{d}s\|_{L^{p}}\lesssim\|f\|_{C_{t}\mathcal{C}^{\alpha}}2^{j\beta}\|\Delta_{j}g(t)\|_{L^{p}}$
since $S_{j-1}f,\Lambda^{\gamma}\widetilde{\Delta}_{j}]$ is spectrally supported in dyadic annuli, by Lemma 2.23 in [BCD11], we obtain the estimate.
For $\partial_{t}(f\prec\!\!\!\prec g)-f\prec\!\!\!\prec\partial_{t}g$, it’s suffice to estimate
@@
Acknowledgments This work was partially supported by the National Natural Science Foundation of China (No.12571261).